OpenStudy (anonymous):

What is the value of c so that y = x^2 + 9x + c is a perfect trinomial? a) 18 b) 9/2 c) 9/4 d) 81/4 I WILL GIVE MEDAL AND BECOME A FAN

OpenStudy (anonymous):

@ravikr.iit

OpenStudy (anonymous):

Is D the correct answer?

OpenStudy (theeric):

You can test it yourself! I don't believe that it is the correct answer...

OpenStudy (theeric):

Or maybe it is...

OpenStudy (theeric):

It might be...

OpenStudy (accessdenied):

A perfect trinomial factors into a squared binomial like this: \( (x + h)^2 \) This expands out to \(x^2 + 2h x + h^2 \) We can then compare coefficients between the trinomials. 2h = 9 h squared will become the necessary addition to make a perfect square trinomial, c = h^2.

OpenStudy (theeric):

Either way, whoever does know for sure should show justification for learning purposes. Thank you @AccessDenied .

OpenStudy (theeric):

So, I agree with both of you now :)

OpenStudy (anonymous):

yeah \[(x+9/2)^{2}\] is perfect trinomial

OpenStudy (anonymous):

Thanks everyone.

OpenStudy (theeric):

I didn't do too much, but you're welcome :)

OpenStudy (anonymous):

I have a few more questions, here is one of them: What are the solutions of the equation? 0=x^2+3x-10 a)x+5,2 b)x=-4,-2 c)x=-5,2 d)x=5,-2

OpenStudy (anonymous):

a)x=5,2* sorry

OpenStudy (accessdenied):

Have you attempted to factor the quadratic? Often this is helpful to split it into single linear equations. A*B = 0 means A=0 or B=0.

OpenStudy (anonymous):

FYI, it's not called a perfect trinomial. It's called a perfect square trinomial.

OpenStudy (anonymous):

I have one more question: A ball is thrown into the air with an upward velocity of 32 feet per second. Its height, h, in feet after t seconds is given the function h(t)=-16t^2+6. How long does it take the ball to reach its maximum height? What is the ball's maximum height? Round to the nearest tenth, if necessary. a)reaches a maximum height of 22 feet after 1.00 seconds. b) reaches a maximum height of 22 feet after 2.00 seconds. c)reaches a maximum height of 44 feet after 2.17 seconds. d)reaches a maximum height of 11 feet after 2.17 seconds. @AccessDenied @ravikr.iit @theEric @ranga can you help?

OpenStudy (anonymous):

at max. hight velocity=o so differentiate the function h(t) and u'll find velocity put is zero and solve for t u'll get the time for max hight. then putt this value in function of hight h(t) u'll get the max hight

OpenStudy (accessdenied):

that would be the right idea, although i suspect this is algebra 2 and the operation of differentiation may not be known. here. instead, there is an alternative using the vertex form or x = -b/2a to find vertex of the parabola, the high point.

OpenStudy (theeric):

But we can just look at it more simply... In case we haven't learned differentiation. What \(t\) will make that function largest? Well, the \(t\) multiplies a negative...... Ooo.... @AccessDenied has a MUCH more relevant idea :)

OpenStudy (theeric):

And it gives the same result! :)

OpenStudy (anonymous):

Okay thanks again(:

OpenStudy (accessdenied):

you're welcome! :)

OpenStudy (theeric):

Take care!

OpenStudy (ranga):

Are you sure h(t) = -16t^2 + 6? Is the "t" term missing?

OpenStudy (theeric):

@ranga it is odd... I just assumed it to be written correctly, although it contradict the wording "thrown into the air." Without that \(t\) term, it is not thrown. Rather, it is dropped from six feet above the ground somewhere very roughly around sea level...

OpenStudy (theeric):

@nichole_rawr_ did you see what ranga posted?